OpenStudy (anonymous):

How to solve for the inverse of this function. f(x)=(x+1)/2

OpenStudy (anonymous):

Let f(x)=y you get y=(x+1)/2

OpenStudy (anonymous):

Then flip flop x and y, then solve for y x=(y+1)/2 2x=y+1 y=2x-1

OpenStudy (anonymous):

thank you!